# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ 6.3 Covariate Shift Adaptation (Density Ratio Estimation)
We consider the covariate shift setting in Section 2. Let $X$ be the source covariate distribution that generates labeled observations $\{(X_{i},Y_{i})\}^{n}_{i=1}$, and let $\widetilde{X}$ be the target covariate distribution that generates unlabeled observations $\{\widetilde{X}_{j}\}^{m}_{j=1}$, independent of the source sample. Let $p_{0}(x)$ and $p_{1}(x)$ be the pdfs of $X$ and $\widetilde{X}$, repsecitively. We assume that $p_{0}(x),p_{1}(x)>0$ for all $x\in{\mathcal{X}}$. The Riesz representer for covariate shift adaptation is the density ratio
@@
Table 2: Experimental results using the synthetic dataset. CR denotes the coverage ratio of 9595% confidence intervals; that is, values close to $0.95$ are better. DM denotes the direct method, which is independent of the generalized Riesz regression; therefore, in theory, the results of the DM estimator should not differ across SQ-Riesz regression (SQ-Riesz) and UKL-Riesz regression (UKL-Riesz). Since we compute the DM estimator when constructing the AIPW estimator in each of SQ-Riesz regression and UKL-Riesz regression, we also report the DM estimator results for reference.
True SQ-Riesz UKL-Riesz
DM IPW AIPW DM IPW AIPW DM IPW AIPW
MSE 0.00 1.10 0.01 0.30 0.59 0.11 0.31 0.36 0.09
CR 1.00 0.92 0.97 0.17 0.97 0.87 0.11 0.92 0.87
@@ 7 Experiments
We assess the performance of our method through simulation studies that evaluate ATE estimation error.
@@ 7.1 Experiments with Synthetic Dataset
We first conduct simulation studies to examine the empirical behavior of our generalized Riesz regression under different choices of Bregman divergence. We focus on three representative losses introduced in Section 3: SQ-Riesz regression and the UKL-Riesz regression. In this experiment, we explicitly use cross fitting in the sense of Assumption 5.4. This setting illustrates how our framework can be combined with high capacity models without relying on the Donsker assumption.
Design and implementation
We consider two covariate dimensions, $K=3$ and $K=10$, and two outcome models. This yields four experimental settings. In all cases, the true ATE is fixed at $\theta_{0}=5.0$. We generate covariates $X_{i}\in{\mathbb{R}}^{K}$ from a multivariate normal distribution $\mathcal{N}(0,I_{K})$, where $I_{K}$ denotes the $K\times K$ identity matrix. The propensity score is defined as $e_{0}(X_{i})=\frac{1}{1+\exp\bigl(-\alpha(X_{i})\bigr)}$, where
$\alpha(X_{i})=\sum_{j=1}^{3}\alpha_{j}X_{i,j}+\sum_{j=1}^{3}\beta_{j}X_{i,j}^{2}+\gamma_{1}X_{i,1}X_{i,2}+\gamma_{2}X_{i,2}X_{i,3}+\gamma_{3}X_{i,1}X_{i,3}.$
@@
The coefficients $\alpha_{j}$, $\beta_{j}$, and $\gamma_{j}$ are independently drawn from $\mathcal{N}(0,0.5)$. Given these propensity scores, the treatment assignment $D_{i}$ is sampled from a Bernoulli distribution with success probability $e_{0}(X_{i})$. The outcome is then generated under two models. In Model 1, we specify
$Y_{i}=\left(X^{\top}_{i}\beta\right)^{2}+1.1+\theta_{0}D_{i}+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$. In Model 2, the outcome is generated as
$Y_{i}=X^{\top}_{i}\beta+\left(X^{\top}_{i}\beta\right)^{2}+3\sin(X_{i,1})+1.1+\theta_{0}D_{i}+\varepsilon_{i}.$
To evaluate efficiency and coverage properties, we construct an oracle benchmark that uses the true nuisance functions. For each replication, we compute infeasible DM, IPW, and AIPW estimators based on the true propensity score and the true conditional expectations of $Y(d)$, and we use their corresponding influence functions to form oracle 9595% confidence intervals. The performance of these oracle estimators is summarized in the “True” columns of Table 2.
For our proposed DBC estimators, we estimate the bias correction term $\alpha_{0}(X)$ using fully connected neural networks with one hidden layer of 100 nodes. For DBC (LS), we estimate $\alpha_{0}$ directly using the squared loss objective associated with Riesz regression. For DBC (KL) and DBC (TL), we model $\alpha_{0}$ by estimating the propensity score and plugging it into the ATE Riesz representer, using the KL divergence loss and the tailored loss introduced in Section 3.4, respectively. The conditional outcome regression $\gamma_{0}(X)$ used for the DM and AIPW estimators is also modeled by a neural network with one hidden layer and 100 nodes. For the MLE baseline, we estimate the propensity score using the same neural network architecture and the Bernoulli log likelihood.
@@
To avoid relying on the Donsker condition, all nuisance functions are estimated with two fold cross fitting. In each replication, we split the sample into two folds, estimate nuisance functions on one fold, evaluate the corresponding scores on the other fold, and then swap the roles of the folds. The final estimators aggregate the two cross fitted scores.
For each loss (LS, KL, TL), we report three estimators:
  • the direct method (DM), which depends only on the outcome regression,
  • the IPW estimator, constructed using the estimated bias correction term,
  • the AIPW estimator, which combines the estimated bias correction term and the outcome regression.
The DM estimator is theoretically independent of the loss used to estimate the bias correction term. In practice, we recompute the DM estimator within each DBC run to construct the AIPW estimator, and we report the resulting DM performance for reference. Small differences among the DM columns therefore reflect only Monte Carlo variation.
We repeat the experiment 100 times. For each method and each estimator (DM, IPW, AIPW), we compute the empirical mean squared error (MSE) of the ATE estimate and the empirical coverage ratio (CR) of the nominal 9595% confidence interval, defined as the fraction of replications in which the interval contains the true effect $\theta_{0}$. The results are summarized in Table 2.
@@ Results
Table 2 reports the MSE and coverage ratio for the oracle estimators (True) and for the three DBC variants. The oracle AIPW estimator achieves a very small MSE (approximately $0.01$) and a coverage ratio close to the nominal level (about $0.97$), as expected. The oracle IPW estimator has a larger MSE (around $1.10$) and slightly conservative coverage (about $0.92$). The oracle DM estimator is unbiased by construction, so its MSE is essentially zero and its coverage ratio is close to one.
For the feasible DBC estimators, the DM columns are nearly identical across DBC (LS), DBC (KL), and DBC (TL), with MSE around $0.30$ and poor coverage (CR between $0.11$ and $0.17$). This behavior reflects the well known fact that the plug in DM estimator is not debiased and is not suitable for inference in this design, even when the outcome model is reasonably flexible.
The IPW estimators based on our estimated bias correction term exhibit substantially reduced MSE relative to the oracle IPW benchmark that uses the true propensity score. This type of “paradox” is reported and analyzed in existing studies, such as Hirano et al. (2003) and Henmi & Eguchi (2004). Under DBC (LS), the IPW MSE is about $0.59$, while DBC (KL) and DBC (TL) further reduce it to approximately $0.41$ and $0.36$, respectively. The coverage ratios for IPW are close to the nominal level for all three losses (around $0.97$ for DBC (LS) and DBC (KL), and about $0.92$ for DBC (TL)). These results indicate that direct estimation of the bias correction term can improve both efficiency and coverage for IPW, and that the KL and tailored losses provide modest gains over the squared loss in this setting.
The AIPW estimators exhibit the best overall performance. All three DBC variants achieve small MSEs, with values around $0.11$ for DBC (LS), $0.08$ for DBC (KL), and $0.09$ for DBC (TL), which are close to the oracle AIPW MSE of $0.01$. The coverage ratios of the AIPW estimators are slightly below the nominal level (between $0.87$ and $0.88$) but still reasonably close, given the moderate number of Monte Carlo replications. The differences among the three losses are minor, with DBC (KL) and DBC (TL) showing a slight advantage in terms of MSE.
Overall, these additional experiments support our theoretical findings. First, they confirm that direct estimation of the bias correction term via Bregman divergence minimization yields ATE estimators that are close to the oracle benchmark when combined with cross fitting. Second, they show that the choice of Bregman divergence (squared loss, KL loss, or tailored loss) has only a modest impact on the AIPW estimator, while the KL and tailored losses can yield small efficiency gains in some cases. Third, they illustrate that our framework can be implemented with flexible neural network models and cross fitting, without relying on the Donsker condition.
@@ 7.2 Experiments with semi synthetic datasets
We next evaluate the proposed estimators on a semi synthetic benchmark based on the Infant Health and Development Program (IHDP) data, following Chernozhukov et al. (2022a). The IHDP was a randomized trial that investigated the effect of an early childhood intervention on subsequent developmental and health outcomes. Following the standard setting “A” implemented in the npci package, we generate 1000 semi synthetic datasets, each consisting of $n=747$ observations with a binary treatment $T$, an outcome $Y$, and $p=25$ continuous and binary covariates $X$. The estimand of interest is the average treatment effect (ATE) of the intervention on $Y$.
For each dataset, we compute three ATE estimators, the direct method (DM), the inverse probability weighting (IPW) estimator, and the augmented IPW (AIPW) estimator. All estimators use our generalized Riesz regression for estimating the Riesz representer.
The nuisance functions are estimated either by a feedforward neural network or by an RKHS regression. The neural network has a single hidden layer with 100 units and is trained for 100 epochs. For the RKHS learner, we use 100 Gaussian basis functions. The bandwidth of the Gaussian kernel and the ridge regularization parameter are chosen by cross validation.
To assess estimation accuracy and uncertainty quantification, we report the mean squared error (MSE) of each ATE estimator and the empirical coverage ratio (CR) of nominal 9595% Wald type confidence intervals across the 1000 replications. Here, CR is defined as the proportion of replications in which the confidence interval contains the true ATE, so values close to $0.95$ indicate well calibrated intervals. The results are summarized in Table 3.
Overall, when neural networks are used for nuisance estimation, the AIPW estimator combined with our DBC schemes achieves substantially smaller MSE than the corresponding DM and IPW estimators, while its CR is close to one, indicating slightly conservative but reliable inference. The DM estimator exhibits noticeable bias and severe undercoverage, and the IPW estimator can be unstable for some DBC variants. When RKHS learners are employed, the IPW estimator performs relatively well in terms of both MSE and CR, whereas the DM and AIPW estimators are more sensitive to the choice of DBC method and can exhibit larger MSE or poor coverage. These findings suggest that, in this IHDP benchmark, DBC based AIPW with neural network nuisance learners provides the most accurate and well calibrated ATE estimates.
@@ Appendix A Related Work
Construction of efficient estimators in semiparametric analysis is a classical problem in statistics, machine learning, economics, epidemiology, and related fields. In this problem, we consider semiparametric models with a low-dimensional parameter of interest and additional nuisance parameters. Our interest lies in obtaining $\sqrt{n}$-consistent and asymptotically normal estimators of the parameter of interest. The difficulty stems from the estimation error of nuisance parameters, whose convergence rates are typically slower than, or at best comparable to, $\sqrt{n}$. Reducing the influence of nuisance estimation error has been investigated in many studies, including Levit (1976), Ibragimov & Khas’minskii (1981), Robinson (1988), Newey (1994), van der Vaart (1998), Bickel et al. (1998), Ai & Chen (2012), and Chernozhukov et al. (2018).
There are three main approaches to constructing efficient estimators: one-step bias correction, estimating equation methods, and targeted maximum likelihood estimation (TMLE) (Schuler et al., 2018; van der Laan, 2006; van der Laan & Rose, 2018). In all of these approaches, the Riesz representer plays a crucial role. It is often called the bias-correction term in one-step bias correction and the clever covariate in TMLE. The terminology “Riesz representer” appears in the double/debiased machine learning formulation, which organizes existing approaches through the lens of Neyman orthogonality (Chernozhukov et al., 2022b).

### v2

@@ 7.3 Covariate Shift Adaptation (Density Ratio Estimation)
We consider the covariate shift setting in Section 2. Let $X$ be the source covariate distribution that generates labeled observations $\{(X_{i},Y_{i})\}^{n}_{i=1}$, and let $\widetilde{X}$ be the target covariate distribution that generates unlabeled observations $\{\widetilde{X}_{j}\}^{m}_{j=1}$, independent of the source sample. Let $p_{0}(x)$ and $p_{1}(x)$ be the pdfs of $X$ and $\widetilde{X}$, respectively. We assume that $p_{0}(x),p_{1}(x)>0$ for all $x\in{\mathcal{X}}$. The Riesz representer for covariate shift adaptation is the density ratio
@@
8 Overfitting Problems
Non-negative Bregman divergence
Telescoping density ratio estimation
$\frac{p_{0}(x)}{p_{m}(x)}=\prod_{k=0}^{m-1}\frac{p_{k}(x)}{p_{k+1}(x)}.$
Table 2: Experimental results using the synthetic dataset. We report the mean squared error (MSE) of the ATE estimator and the empirical coverage ratio (CR) of nominal 9595% Wald-type confidence intervals over 100 Monte Carlo replications. The column block “True” uses oracle nuisance functions (the true propensity score and true outcome regressions) and is therefore infeasible. “SQ-Riesz” and “UKL-Riesz” estimate the ATE Riesz representer by generalized Riesz regression with the squared-loss and unnormalized-KL objectives, respectively; “BKL-Riesz = MLE” corresponds to estimating the propensity score by BKL objective (logistic MLE) and plugging it into the ATE Riesz representer. For SQ-Riesz we compare two link specifications (Linear and Logit). For UKL-Riesz we compare two feature sets: ${\bm{\phi}}(Z)$ uses only covariates $Z$, while ${\bm{\phi}}(X)$ uses the full regressor $X=(D,Z)$ (allowing treatment-dependent features). For each Riesz-representer fit we report three estimators: the direct method (DM), inverse probability weighting (IPW), and augmented IPW (AIPW). DM depends only on the outcome regression; small differences across DM columns arise from recomputing the outcome regression within each run of the AIPW construction. Values of CR close to $0.95$ indicate well-calibrated uncertainty quantification.
True SQ-Riesz (Linear) SQ-Riesz (Logit) UKL-Riesz (${\bm{\phi}}(Z)$) UKL-Riesz (${\bm{\phi}}(X)$) BKL-Riesz (MLE)
DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW
MSE 0.00 1.44 0.01 0.39 0.49 0.19 0.39 1.38 0.08 0.38 1.50 0.10 0.40 1.52 0.10 0.39 3.79 0.23
CR 1.00 0.84 0.98 0.06 0.98 0.87 0.12 0.80 0.89 0.08 0.73 0.77 0.06 0.68 0.81 0.06 0.32 0.60
@@ 9 Experiments
We evaluate generalized Riesz regression as a building block for debiased machine learning, focusing on average treatment effect (ATE) estimation. Across all experiments, we compare three ways of estimating the ATE Riesz representer (bias-correction term) introduced in Section 3: SQ-Riesz (squared-loss objective), UKL-Riesz (unnormalized-KL objective), and BKL-Riesz (binary-KL objective). In the ATE setting, BKL-Riesz coincides with estimating the propensity score by Bernoulli likelihood (logistic MLE) and then plugging it into the closed-form ATE Riesz representer; we therefore refer to it as “BKL-Riesz = MLE.”
Given an estimate of the outcome regression $\widehat{\gamma}$ and an estimate of the Riesz representer $\widehat{\alpha}$, we report three ATE estimators:
  • DM: the plug-in direct method based only on $\widehat{\gamma}$,
  • IPW: the IPW estimator based only on $\widehat{\alpha}$,
  • AIPW: the Neyman-orthogonal (doubly robust) estimator combining $\widehat{\gamma}$ and $\widehat{\alpha}$ as in Section 2.
We quantify accuracy by the mean squared error (MSE) of the ATE estimate and quantify uncertainty by the empirical coverage ratio (CR) of nominal $95$
@@ 9.1 Experiments with synthetic dataset
Design
The covariates are three-dimensional, $K=3$, and we fix the sample size at $n=3000$. In each Monte Carlo replication, we generate covariates $X_{i}\in{\mathbb{R}}^{3}$ from a multivariate normal distribution $\mathcal{N}(0,I_{3})$ and construct a nonlinear propensity score model with polynomial and interaction terms as
$e_{0}(X_{i})=\frac{1}{1+\exp\bigl(-h(X_{i})\bigr)},$
where
$h(X_{i})=\sum_{j=1}^{3}a_{j}X_{i,j}+\sum_{j=1}^{3}b_{j}X_{i,j}^{2}+c_{1}X_{i,1}X_{i,2}+c_{2}X_{i,2}X_{i,3}+c_{3}X_{i,1}X_{i,3}.$
@@
The coefficients $a_{j}$, $b_{j}$, and $c_{j}$ are independently drawn from $\mathcal{N}(0,0.5)$. Given these propensity scores, the treatment assignment $D_{i}$ is sampled accordingly. We then generate the outcome as
$Y_{i}=1.0+\left(\sum^{3}_{j=1}X_{i,j}\widetilde{a}_{j}\right)^{2}+1/\left(1+\exp\left(-\left(\sum^{3}_{j=1}X^{2}_{i,j}\widetilde{b}_{j}\right)\right)\right)+\tau_{0}D_{i}+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ and $\tau_{0}=5.0$.
Estimators and implementation
We estimate the Riesz representer using the following variants, matched to Table 2.
  • SQ-Riesz (Linear) and SQ-Riesz (Logit): squared-loss generalized Riesz regression with two different link specifications for the Riesz-representer model.
  • UKL-Riesz (${\bm{\phi}}(Z)$) and UKL-Riesz (${\bm{\phi}}(X)$): UKL generalized Riesz regression with a log-type link, comparing two feature sets. Here $X=(D,Z)$ and ${\bm{\phi}}(Z)$ uses only $Z$, while ${\bm{\phi}}(X)$ uses the full regressor (allowing treatment-dependent features).
  • BKL-Riesz (MLE): propensity-score MLE (Bernoulli likelihood) followed by plugging $\widehat{e}(Z)$ into the ATE Riesz representer.
@@
For the Riesz representer and regression models, we separately use a neural network with one hidden layer consisting of $100$ nodes. To avoid relying on the Donsker condition, we estimate all nuisance functions using two-fold cross fitting. In each replication, we split the sample into two folds, estimate the nuisance functions on one fold, evaluate the corresponding scores on the other fold, and then swap the roles of the folds. The final estimators aggregate the two cross-fitted scores.
This experiment does not guarantee automatic Neyman orthogonalization, since we use cross fitting and do not use the same basis functions for outcome modeling. However, this implementation is standard in debiased machine learning; therefore, we adopt it.
We repeat the simulation 100 times. The “True” columns in Table 2 report infeasible oracle performance using the true nuisance functions.
@@ Results
Table 2 highlights three robust patterns. First, oracle baselines separate estimation error from intrinsic variance. The oracle AIPW estimator is close to the efficiency benchmark (MSE =0.01=0.01) and achieves near-nominal coverage (CR =0.98=0.98). In contrast, even with the true propensity score, oracle IPW remains noisy (MSE =1.44=1.44) and undercovers (CR =0.84=0.84), reflecting the well-known finite-sample instability of pure weighting in challenging overlap regimes.
Second, the plug-in DM estimator is not reliable for inference in this design. Across feasible implementations, DM has moderate MSE (about 0.380.380.400.40) but extremely poor coverage (CR =0.06=0.060.120.12). This indicates that the outcome regression learner, while not catastrophically inaccurate in MSE, does not deliver a reliable uncertainty estimate when used without orthogonalization, and the resulting Wald intervals are severely miscalibrated.
Third, how we fit the Riesz representer matters substantially for IPW, and AIPW mitigates (but does not eliminate) this sensitivity. For IPW, SQ-Riesz (Linear) is the best-performing option in Table 2 (MSE =0.49=0.49) and yields near-nominal coverage (CR =0.98=0.98). In contrast, IPW based on UKL-Riesz has larger MSE (about $1.50$) and noticeably lower coverage (CR =0.68=0.680.730.73), while BKL-Riesz (= MLE) performs worst (MSE =3.79=3.79, CR =0.32=0.32), consistent with propensity-score MLE producing more extreme effective weights in this design.
The AIPW estimator is uniformly more stable than IPW and DM in terms of MSE, but calibration still depends on the Riesz-representer fit. SQ-Riesz (Logit) attains the best AIPW MSE (0.08) with CR 0.89. UKL-Riesz achieves similarly small AIPW MSE (0.10) but exhibits undercoverage (CR =0.77=0.770.810.81). BKL-Riesz (= MLE) improves substantially over its IPW counterpart (AIPW MSE =0.23=0.23), yet its coverage remains poor (CR =0.60=0.60). Overall, directly fitting the Riesz representer via generalized Riesz regression can materially improve finite-sample performance relative to the MLE plug-in baseline, and the combination of objective and link specification plays a first-order role, especially for IPW and for the calibration of AIPW intervals.
@@ 9.2 Experiments with semi synthetic datasets
We next evaluate the same family of estimators on the semi-synthetic IHDP benchmark, following Chernozhukov et al. (2022a). We use the standard setting “A” in the npci package and generate 1000 replications. Each replication contains $n=747$ observations with a binary treatment, an outcome, and $p=25$ covariates. The estimand is the ATE.
We report DM, IPW, and AIPW for each Riesz-representer estimator (SQ-Riesz, UKL-Riesz, and BKL-Riesz (= MLE)). We consider two nuisance-learner families:
  • a feedforward neural network with one hidden layer of 100 units,
  • an RKHS learner with 100 Gaussian basis functions (with tuning by cross validation).
For each configuration, we compute the MSE of the ATE estimate and the empirical coverage ratio (CR) of nominal 9595% Wald-type confidence intervals across the 1000 replications; CR close to $0.95$ indicates well-calibrated uncertainty quantification. Results appear in Table 3.
Two findings stand out. With neural networks, AIPW is consistently accurate (MSE around 0.310.310.430.43) and well calibrated for UKL-Riesz and BKL-Riesz (CR =0.94=0.94 and $0.90$), while SQ-Riesz yields overly conservative AIPW intervals (CR =1.00=1.00). In contrast, DM has very low coverage (CR near zero) and IPW exhibits large error, especially for SQ-Riesz (MSE =6.82=6.82), reinforcing that orthogonalization is essential in this benchmark.
With RKHS, performance becomes much more sensitive to the particular objective: SQ-Riesz deteriorates sharply (MSE around $20$ with CR =0=0 for both DM and AIPW), whereas UKL-Riesz and BKL-Riesz remain substantially more stable. In particular, UKL-Riesz attains strong IPW calibration under RKHS (CR =0.93=0.93) with comparatively low MSE (1.78), while BKL-Riesz provides a competitive alternative (IPW MSE =1.22=1.22 with CR =0.81=0.81). These results underscore that, in finite samples, the interaction between the Riesz-representer objective and the nuisance-function learner can be decisive, and that UKL-type objectives can offer noticeably more robust behavior than squared-loss fitting in this semi-synthetic setting.
@@ Appendix A Related Work
A.1 Asymptotically Efficient Estimation
Early history.
The construction of asymptotically efficient estimators is a classical problem in statistics, machine learning, economics, epidemiology, and related fields. In this problem, we consider semiparametric models with a low-dimensional parameter of interest and additional nuisance parameters. Our interest lies in obtaining $\sqrt{n}$-consistent and asymptotically normal estimators of the parameter of interest. The difficulty stems from the estimation error of nuisance parameters, whose convergence rates are typically slower than, or at best comparable to, $\sqrt{n}$. Reducing the influence of nuisance estimation error has been investigated in many studies, including Levit (1976), Ibragimov & Khas’minskii (1981), Pfanzagl & Wefelmeyer (1982), Klaassen (1987), Robinson (1988), Newey (1994), van der Vaart (1998), Bickel et al. (1998), Ai & Chen (2012), and Chernozhukov et al. (2018).
In the construction of asymptotically efficient estimators, we aim to develop estimators that are $\sqrt{n}$-consistent and asymptotically normal, with asymptotic variances that attain the asymptotic efficiency bounds. Asymptotic efficiency bounds are called H’ajek–Le Cam efficiency bounds, or semiparametric efficiency bounds when we consider semiparametric models (Hájek, 1970; Le Cam, 1972, 1986). They share the same motivation as the Cram’er–Rao lower bound. While the Cram’er–Rao lower bound provides a lower bound for unbiased estimators, asymptotic efficiency bounds provide lower bounds for asymptotically unbiased estimators, called regular estimators. It is known that efficient estimators are regular and asymptotically linear (RAL) with the efficient influence function. Therefore, the construction of asymptotically efficient estimators is equivalent to the construction of RAL estimators (van der Vaart, 1998).
There are three main approaches to constructing efficient estimators, one-step bias correction, estimating equation methods, and TMLE (Schuler et al., 2018; van der Vaart, 2002; van der Laan, 2006; van der Laan & Rose, 2018). In many cases, these approaches yield estimators that are asymptotically equivalent. However, their finite sample behavior may differ. Another related line of work is post-selection inference with high-dimensional control variables (Belloni et al., 2011, 2014, 2016).
Debiased machine learning.
Chernozhukov et al. (2018) proposes debiased machine learning as a framework for constructing asymptotically efficient estimators. This framework formulates the problem through Neyman orthogonality, which is closely connected to the semiparametric efficient influence function (Newey, 1994).
Riesz representer.
The Riesz representer appears as part of the Neyman orthogonal score. It is often called the bias-correction term in one-step bias correction and the clever covariate in TMLE. The terminology “Riesz representer” appears in the debiased machine learning formulation based on Neyman orthogonality (Chernozhukov et al., 2022b).
A.2 ATE Estimation
Randomized controlled trials are the gold standard for causal inference, where treatments are assigned while maintaining balance between treatment groups. However, they are not always feasible, and we aim to estimate causal effects from observational data, where imbalance between treatment groups often arises. To correct this imbalance, propensity scores or balancing weights have been proposed.